OpenStudy (anonymous):

how do I solve x/3=-8

OpenStudy (anonymous):

Solve for x: x/3 = -8 Solve for x. Multiply both sides of x/3 = -8 by 3: (3 x)/(3) = 3 (-8) Cancel common terms in the numerator and denominator of (3 x)/3. (3 x)/3 = 3/3×x = x: x = 3 (-8) Multiply 3 and -8 together. 3 (-8) = -24: Answer: | | x = -24

OpenStudy (anonymous):

Just multiply the denominator by -8